# ManiBox: Enhancing Spatial Grasping Generalization via Scalable Simulation Data Generation

### v1

@@ Abstract
Learning a precise robotic grasping policy is crucial for embodied agents operating in complex real-world manipulation tasks. Despite significant advancements, most models still struggle with accurate spatial positioning of objects to be grasped. We first show that this spatial generalization challenge stems primarily from the extensive data requirements for adequate spatial understanding. However, collecting such data with real robots is prohibitively expensive, and relying on simulation data often leads to visual generalization gaps upon deployment. To overcome these challenges, we then focus on state-based policy generalization and present ManiBox, a novel bounding-box-guided manipulation method built on a simulation-based teacher-student framework. The teacher policy efficiently generates scalable simulation data using bounding boxes, which are proven to uniquely determine the objects’ spatial positions. The student policy then utilizes these low-dimensional spatial states to enable zero-shot transfer to real robots. Through comprehensive evaluations in simulated and real-world environments, ManiBox demonstrates a marked improvement in spatial grasping generalization and adaptability to diverse objects and backgrounds. Further, our empirical study into scaling laws for policy performance indicates that spatial volume generalization scales positively with data volume. For a certain level of spatial volume, the success rate of grasping empirically follows Michaelis-Menten kinetics relative to data volume, showing a saturation effect as data increases. Our videos and code are available in the project page.
Figure 1: Overview. We introduce ManiBox, a bbox-guided manipulation method using a teacher-student framework to enhance spatial generalization in grasping policies. We reveal that spatial volume generalization scales positively with data volume, with grasping success following Michaelis-Menten kinetics relative to data volume for specified spatial volumes. Extensive real-world tests show ManiBox’s robust adaptability to varied spatial positions, objects, and backgrounds.
@@ 1 Introduction
Recent works have made notable progress in robotic manipulation in multi-task and multi-environment scenarios, which is largely driven by end-to-end training methodologies (Ahn et al., 2022; Shridhar et al., 2022; 2023; Zitkovich et al., 2023). Notably, some large embodied foundation models for manipulation, such as those fine-tuned from Vision-Language Models (VLMs) for multi-task robot manipulation tasks (Brohan et al., 2022; Kim et al., 2024), have made significant strides. However, these large manipulation models derived from VLMs often lack sufficient spatial awareness and reasoning capabilities, posing substantial challenges for large-scale spatial generalization in real-world applications (Cheng et al., 2024a). Our analysis in Lemma 1 suggests that achieving spatial generalization over larger volumes requires substantially more data, a relationship we further empirically verify to follow a positive scaling law (see Figure 4). For example, training a model to grasp at a fixed point may require only 50 to 600 trajectory datasets. However, when generalizing the model to operate within a spatial volume of $34,400cm^{3}$ ($41cm\times 30cm\times 28cm$), which approximates the maximum reach of our robotic arm, the data requirement increases by roughly 40-fold, demanding tens of thousands of trajectories
@@
While perception models are essential for object and environmental recognition, the true effectiveness of a robot in dynamic settings hinges on the robustness of its policy execution. A well-generalized policy empowers the robot to autonomously adjust its actions in real-time, effectively compensating for visual inaccuracies and adapting to changes in object positions or environmental dynamics. Such adaptability is crucial for achieving task success and maintaining system stability, especially in the face of imperfect or noisy perception data. To further enhance the robot’s ability to manage unpredictable scenarios and ensure reliable real-world performance, we shift our focus to state-based policy generalization instead of relying solely on vision-based methods. Our Lemma 2 demonstrates that bounding boxes, when captured from multiple cameras, effectively encapsulate the 3D structure of convex objects, providing an optimal low-dimensional state for policy generalization. Building on this foundation, we introduce a novel bounding-box-guided methodology ManiBox to enhance spatial generalization and adaptability.
@@ 3.2 Teacher Policy Training
However, due to challenges such as the high dimensionality of the state space and the large exploration space caused by the multi-stage reward function during the training of the teacher policy, the number of samples required for RL training is significantly large. Therefore, in addition to optimizing the policy using the Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm, we also make adjustments to the RL environment to reduce the exploration space and make it easier for the policy $\pi_{\beta}$ to find the point with the highest return (More details are in Appendix C.1):
@@ 3.3 Simulation Data Generation
This lemma illustrates that 3D position information of a sphere can be implied by utilizing multi-view bounding boxes, indicating our ManiBox can capture complete spatial position information of target objects for real-world control policies.
@@ 3.4 Student Policy Distillation
where $\mathrm{LSTM}(\cdot),\mathrm{gelu}(\cdot),\mathrm{drop}(\cdot),\mathrm{FC_{\cdot}}(\cdot)$ represent the Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997), Gaussian Error Linear Unit (GELU) activation (Hendrycks & Gimpel, 2016), dropout (Srivastava et al., 2014), and Fully Connected layer (FC) respectively.
This combination of random mask and trajectory modeling significantly contributes to the robustness and generalization of our student policy in deployment, conquering the challenge of partial observations and unseen real-world variations.
@@ 4.2 Spatial Scaling Law with Data
Figure 3: The scaling relationship between spatial generalization and data amounts under different spatial ranges in the grasping task. The data volume is the number of trajectories used for training the student policy. The estimated point with an 80% success rate is annotated additionally.
@@
Figure 4: The relationship between spatial volume and data amounts needed to reach 80% grasping success rate. The fitted curve represents a power function of $y=1507\cdot x^{0.28}$.
@@
Furthermore, we identify the data volume required to achieve an 80% success rate for each spatial range (as shown in Fig. 3) and find that the volume of the spatial range and the required data follow a power rate relationship (as shown in Fig. 4). The positive correlation between data volume and spatial volume provides guidance for data-driven spatial generalization: The more data we have, the higher the robot’s manipulation success rate. As the spatial range increases, more data is required, and the data volume needed grows with the space volume as a power rate relationship. In fact, for the same task, achieving high generalization success over a $41cm*30cm*28cm=34440cm^{3}$ range may require at least 20,000 trajectory data from different spatial positions. Such large-scale data collection would be expensive in the real world, making simulator-generated data a more affordable and efficient option. This further emphasizes the significance of simulation data.
@@ 4.3 Real-world Results
To validate the generalization ability of our method and policy in the real world, we compare the success rate of our policy with the vision-based ACT. In practice, we deploy models trained on simulation data onto a real robot and test their performance on grasping tasks across multiple arbitrarily chosen positions, eight objects(with some of them out-of-distribution), five backgrounds, and three unseen settings as illustrated in Figure 5. More images and videos of these tests are in the Appendix E.3 and supplementary materials.
@@
To justify the efficacy of our method, we assess the generalization capability of our student policy across four spatial ranges and a fixed point. For each experiment, we randomly place an apple, different from its counterpart in the simulator in both shape and material, on the table at arbitrarily chosen heights within the given range(as illustrated in the first three rows of Fig. 5). Alongside the difference in form and appearance, the apple presents considerable unpredictability to the visual detection model due to factors including background, lighting, target object, etc. Consequently, the bounding box varies in size or even disappears at times, emphasizing the need for robustness in the policy. As shown in Tab. 1, our policy generalizes well to objects falling in different spatial volumes, achieving a success rate between 70% and 100%. In contrast, vision-based ACT, which trains the student policy with the vision input ACT (Zhao et al., 2023), experiences a visible drop on success rate when spatial range scales up, despite its considerable success rate of 70% on a fixed point. The wanting performance of the baseline further demonstrates the significance and necessity of exploring spatial generalization. The results demonstrate that our method harbors strong generalization when addressing uncertainties arising from real-world objects, variations in spatial positions, and visual detection interferences.
@@
In addition, to justify the object generalization ability of our method, we test our model across eight objects including original apple, unseen apple, yellow pear, orange, steel cup, tin can, glass beaker, and plastic bottle at arbitrarily selected positions. For each unseen object, feeding the corresponding name as a supplement into the open-vocabulary visual detection model is enough for the policy to grasp the target object. It is worth noting that two of the objects chosen are out-of-distribution (their bounding boxes are unseen in the simulation data) in an attempt to further illustrate the object generalization ability of our method. This certain level of generalization observed, even when there is significant variation in the bounding box, aligns with our Lemma 2, which proves that the 3D position information of the object is complete in the dataset.
@@
Utilizing the strong open-vocab detection ability of YOLO-World, ManiBox can handle various backgrounds and vision disruptions without intentionally varied visual domain randomization. To verify this ability, we experiment with tables of three different shapes and tablecloths in two colors under the minimal setting of a table, an object, and a robot. In addition, we alter the lighting conditions by employing a single stationary point light source and a single flickering point light, as well as playing videos in the background as distractions. Furthermore, we construct two more complex scenarios, one representing a multi-object environment and the other replicating a typical daily workplace desktop. The results demonstrate the strong background generalization ability of our policy. Even though the visual detection model occasionally makes mistakes, the generalization of our policy compensates for those errors. This can be attributed to our policy’s endeavor to integrate historical information, multi-camera data, and random masking to increase the visual model’s tolerance for errors.
@@ 5 Conclusion
In this work, we present ManiBox, a novel bbox-guided manipulation method, using a simulation-based teacher-student framework to enhance the spatial generalization of grasping policies. To address the challenge of the generalization gap in vision-based models from simulation to real-world deployment, we propose a generalizable, state-based student policy that is innovatively equipped with bbox positions. This policy is trained on scalable simulation data generated by the teacher policy. Extensive experiments in both simulated and real-world environments demonstrate that ManiBox significantly outperforms existing methods in adapting to diverse spatial positions, object configurations, and background variations. Moreover, our theoretical and empirical analyses reveal that the generalization of spatial volume scales positively with the data volume, and the success rate follows Michaelis-Menten kinetics for specified spatial volumes. These findings pave the way for the development of robust spatial generalization across a wide range of embodied systems.
@@ Appendix B Ablation Study
Training with Only Key Steps. To assess the impact of temporal information on policy effectiveness, we conduct an experiment where the model is trained using only key steps from the trajectory. The specific key steps selected were [0, 18, 20, 22, 70], chosen to represent pivotal moments in the trajectory for the grasping task: 0 corresponds to the initial state before the gripper starts approaching the apple, 70 represents the final state when the robotic arm retracts, and steps 18 to 22 denote intermediate states where the gripper makes contact with the apple and closes around it.
@@
Different Mask Ratios during Inference. During the inference phase, visual models such as YOLO-World may not always successfully identify objects, leading to significant uncertainty in visual input. To test our policy’s robustness against this visual uncertainty, we introduce masks that randomly obscured bounding box input at varying ratios (0.0 to 1.0).
@@
As shown in Fig. 6, the results reveal that the policy maintains a high success rate and generalization capability when the mask ratio is below 0.4. As the mask ratio exceeds 0.4, the success rate significantly declines, leading to consistent policy failures. This trend reveals that our approach focusing on policy generalizability can resist visual uncertainty up to 40% of the masking ratio. That is, even if 40% of the visual information is lost, the policy still has good spatial generalizability. We fitted the data points with a Gaussian curve to represent this trend, defined by the equation:
@@
This experiment validates that our policy can withstand moderate visual uncertainties during inference, maintaining robust performance. However, the steep decline in success rate for higher mask ratios highlights the importance of reliable visual input for effective spatial reasoning and task execution.
Random Noise in Bounding Boxes. To further examine the policy’s robustness against visual uncertainty, we introduce random noise to the bounding boxes during inference. We tested noise ratios at increments of 0.01, 0.02, 0.05, 0.1, 0.2, and 0.5. Fig. 7 illustrates the success rate as a function of noise ratio, where a monotonic decay fit highlights the policy’s performance trend. The results show that the policy maintains a high success rate at low noise ratios but experienced a significant decrease as noise levels increased. Since our normalized bounding box value ranges from 0 to 1, the policy still achieves a high success rate with a visual detection error of 0.05, demonstrating the robustness of the policy generalization to cope with visual uncertainty and visual sim2real gap.

### v2

@@ Abstract
Learning a precise robotic grasping policy is crucial for embodied agents operating in complex real-world manipulation tasks. Despite significant advancements, most models still struggle with accurate spatial positioning of objects to be grasped. We first show that this spatial generalization challenge stems primarily from the extensive data requirements for adequate spatial understanding. However, collecting such data with real robots is prohibitively expensive, and relying on simulation data often leads to visual generalization gaps upon deployment. To overcome these challenges, we then focus on state-based policy generalization and present ManiBox, a novel bounding-box-guided manipulation method built on a simulation-based teacher-student framework. The teacher policy efficiently generates scalable simulation data using bounding boxes, which are proven to uniquely determine the objects’ spatial positions. The student policy then utilizes these low-dimensional spatial states to enable zero-shot transfer to real robots. Through comprehensive evaluations in simulated and real-world environments, ManiBox demonstrates a marked improvement in spatial grasping generalization and adaptability to diverse objects and backgrounds. Further, our empirical study into scaling laws for policy performance indicates that spatial volume generalization scales with data volume in a power law. For a certain level of spatial volume, the success rate of grasping empirically follows Michaelis-Menten kinetics relative to data volume, showing a saturation effect as data increases. Our videos and code are available in the project page.
Figure 1: Overview. We introduce ManiBox, a bbox-guided manipulation method using a teacher-student framework to enhance spatial generalization in grasping policies. We reveal that spatial volume generalization scales with data volume in power law, with grasping success following Michaelis-Menten kinetics relative to data volume for specified spatial volumes. Extensive real-world tests show ManiBox’s robust adaptability to varied spatial positions, objects, and backgrounds.
@@ 1 Introduction
Recent works have made notable progress in robotic manipulation in multi-task and multi-environment scenarios, which is largely driven by end-to-end training methodologies (Ahn et al., 2022; Shridhar et al., 2022; 2023; Zitkovich et al., 2023). Notably, some large embodied foundation models for manipulation, such as those fine-tuned from Vision-Language Models (VLMs) for multi-task robot manipulation tasks (Brohan et al., 2022; Kim et al., 2024), have made significant strides. However, these large manipulation models derived from VLMs often lack sufficient spatial awareness and reasoning capabilities, posing substantial challenges for large-scale spatial generalization in real-world applications (Cheng et al., 2024a). Our analysis in Lemma 1 suggests that achieving spatial generalization over larger volumes requires substantially more data, a relationship we further empirically verify to follow a power scaling law (see Figure 4). For example, training a model to grasp at a fixed point may require only 50 to 600 trajectory datasets. However, when generalizing the model to operate within a spatial volume of $34,400cm^{3}$ ($41cm\times 30cm\times 28cm$), which approximates the maximum reach of our robotic arm, the data requirement increases by roughly 40-fold, demanding tens of thousands of trajectories
@@
While perception models are essential for object and environmental recognition, the true effectiveness of a robot in dynamic settings hinges on the robustness of its policy execution. A well-generalized policy empowers the robot to autonomously adjust its actions in real-time, effectively compensating for visual inaccuracies and adapting to changes in object positions or environmental dynamics. Such adaptability is crucial for achieving task success and maintaining system stability, especially in the face of imperfect or noisy perception data. To further enhance the robot’s ability to manage unpredictable scenarios (e.g. Sec. B.2) and ensure reliable real-world performance, we shift our focus to state-based policy generalization instead of relying solely on vision-based methods. Our Lemma 2 demonstrates that bounding boxes, when captured from multiple cameras, effectively encapsulate the 3D structure of convex objects, providing an optimal low-dimensional state for policy generalization. Building on this foundation, we introduce a novel bounding-box-guided methodology ManiBox to enhance spatial generalization and adaptability.
@@ 3.2 Teacher Policy Training
The advantages of state-based teacher policies lie in their lower-dimensional states, enabling more efficient learning, reduced exploration space through privileged information, and support for larger parallel environments. To leverage these benefits, in addition to optimizing the policy using the Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm, we make adjustments to the RL environment to reduce the exploration space and make it easier for the policy $\pi_{\beta}$ to find the point with the highest return. Moreover, to support follow-up Sim2Real and spatial generalization, we carefully design the environment below (More details are in Appendix E.1):
@@ 3.3 Simulation Data Generation
This lemma illustrates that 3D position information of a sphere can be implied by utilizing multi-view bounding boxes (Hartley & Zisserman, 2003), indicating our ManiBox can capture complete spatial position information of target objects for real-world control policies.
@@ 3.4 Student Policy Distillation
where ${\bm{s}}^{\prime}_{t}$ and ${\bm{h}}_{t}$ denotes the input state and the hidden state of the LSTM at time $t$, ${\bm{x}}_{t}$ is the intermediate output of the LSTM, and ${\bm{a}}_{t}$ is the action vector output by the policy. Furthermore, $\mathrm{LSTM}(\cdot),\mathrm{gelu}(\cdot),\mathrm{drop}(\cdot),\mathrm{FC_{\cdot}}(\cdot)$ represent the Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997), Gaussian Error Linear Unit (GELU) activation (Hendrycks & Gimpel, 2016), dropout (Srivastava et al., 2014), and Fully Connected layer (FC) respectively.
This combination of random mask and trajectory modeling significantly contributes to the robustness and generalization of our student policy in deployment, conquering the challenge of partial observations and unseen real-world variations. (More Sim2Real techniques in Appendix E.4)
@@ 4.2 Spatial Scaling Law with Data
Figure 3: The scaling relationship between spatial generalization and data volume in the grasping task, measured under different spatial ranges. Data volume represents the number of trajectories used to train the student policy. The estimated 80% success point is marked, with blue points showing the average success rate across three seeds and error bars indicating standard deviation.
@@
Figure 4: The relationship between spatial volume and data amounts needed to reach 80% grasping success rate. The fitted curve represents a power function $y=640.32\cdot x^{0.35}$.
@@
Furthermore, we identify the data volume required to achieve an 80% success rate for each spatial range (as shown in Fig. 3) and find that the volume of the spatial range and the required data follow a power rate relationship (as shown in Fig. 4). The power-law correlation between data volume and spatial volume provides guidance for data-driven spatial generalization: The more data we have, the higher the robot’s manipulation success rate. As the spatial range increases, more data is required, and the data volume needed grows with the space volume as a power rate relationship. In fact, for the same task, achieving high generalization success over a $41cm*30cm*28cm=34440cm^{3}$ range may require at least 20,000 trajectory data from different spatial positions. Such large-scale data collection would be expensive in the real world, making simulator-generated data a more affordable and efficient option. This further emphasizes the significance of simulation data.
@@ 4.3 Real-world Results
To validate the generalization ability of our method and policy in the real world, we compare the success rate of our policy with the vision-based ACT. In practice, we deploy models trained on simulation data onto a real robot and test their performance on grasping tasks across multiple arbitrarily chosen positions, eight objects(with some of them out-of-distribution), five backgrounds, and three complex settings as illustrated in Figure 5. More tasks, pictures, and videos of the experiment are available in the Appendix C and supplementary materials. More experiment details of these tests are in the Appendix D.3.
@@
To justify the efficacy of our method, we assess the generalization capability of our student policy across four spatial ranges and a fixed point. For each experiment, we randomly place an apple, different from its counterpart in the simulator in both shape and material, on the table at arbitrarily chosen heights within the given range(as illustrated in the first three rows of Fig. 5). Alongside the difference in form and appearance, the apple presents considerable unpredictability to the visual detection model due to factors including background, lighting, target object, etc. Consequently, the bounding box varies in size or even disappears at times, emphasizing the need for robustness in the policy. As shown in Tab. 1, our policy generalizes well to objects falling in different spatial volumes, achieving a success rate between 70% and 100%. In contrast, vision-based ACT, which trains the student policy with the vision input ACT (Zhao et al., 2023) on video-action simulation data, experiences a visible drop on success rate when spatial range scales up, despite its considerable success rate of 70% on a fixed point. The wanting performance of the baseline further demonstrates the significance and necessity of exploring spatial generalization. The results demonstrate that our method harbors strong generalization when addressing uncertainties arising from real-world objects, variations in spatial positions, and visual detection interferences.
@@
In addition, to justify the object generalization ability of our method, we test our model across eight objects including original apple, unseen apple, yellow pear, orange, steel cup, tin can, glass beaker, and plastic bottle at arbitrarily selected positions. For each unseen object, feeding the corresponding name as a supplement into the open-vocabulary visual detection model is enough for the policy to grasp the target object. Tab. 5 shows detailed object size information for training and inference. It is worth noting that five of the selected objects are out-of-distribution in size (their bounding box sizes are unseen in the simulation data) in an attempt to further illustrate the object generalization ability of our method. This certain level of generalization observed, even when there is significant variation in the bounding box, aligns with our Lemma 2, which proves that the 3D position information of the object is complete in the dataset.
@@
Utilizing the open-vocab detection ability of YOLO-World, ManiBox can handle various backgrounds and vision disruptions without intentionally varied visual domain randomization. To verify this ability, we experiment with tables of three different shapes and tablecloths in two colors under the minimal setting of a table, an object, and a robot. In addition, we alter the lighting conditions by employing a single stationary point light source and a single flickering point light, as well as playing videos in the background as distractions. Furthermore, we construct two more complex scenarios, one representing a multi-object environment and the other replicating a typical daily workplace desktop. The results demonstrate the strong background generalization ability of our policy. Even though the visual detection model occasionally makes mistakes, the generalization of our policy compensates for those errors. This can be attributed to our policy’s endeavor to integrate historical information and random masking to increase the visual model’s tolerance for errors. Arguments for them can be found in Appendix B.
@@ 5 Conclusion
In this work, we present ManiBox, a novel bbox-guided manipulation method, using a simulation-based teacher-student framework to enhance the spatial generalization of grasping policies. To address the challenge of the generalization gap in vision-based models from simulation to real-world deployment, we propose a generalizable, state-based student policy that is innovatively equipped with bbox positions. This policy is trained on scalable simulation data generated by the teacher policy. Extensive experiments in both simulated and real-world environments demonstrate that ManiBox significantly outperforms existing methods in adapting to diverse spatial positions, object configurations, and background variations. Moreover, our theoretical and empirical analyses reveal that the generalization of spatial volume scales with the data volume in power law, and the success rate follows Michaelis-Menten kinetics for specified spatial volumes. These findings pave the way for the development of robust spatial generalization across a wide range of embodied systems.
@@ Appendix B Ablation Study
B.1 Training with Only Key Steps.
To assess the impact of temporal information on policy effectiveness, we conduct an experiment where the model is trained using only key steps from the trajectory. The specific key steps selected were [0, 18, 20, 22, 70], chosen to represent pivotal moments in the trajectory for the grasping task: 0 corresponds to the initial state before the gripper starts approaching the apple, 70 represents the final state when the robotic arm retracts, and steps 18 to 22 denote intermediate states where the gripper makes contact with the apple and closes around it.
@@
B.2 Robustness against Detection Failure during Inference.
During the inference phase, visual models such as YOLO-World may not always successfully identify objects, leading to significant uncertainty in visual input.
Our policy still works well in situations where the visual model fails to detect.
For example, visual models may fail to detect target objects when the robotic arm approaches the target, under dim lighting conditions, or when there is partial occlusion of the target by the background (see Figures 6 and 7). Especially, when using flickering and swaying point light sources in dimly lit environments, the detection by visual models is particularly unstable. Such failures may occur at any timestep during manipulation and may persist for several timesteps. Moreover, these failures are unstable and unpredictable. For the same object placement, the occurrence of detection failures during manipulation does not consistently happen or not happen at the same timestep.
Despite all those failures and instabilities in detecting bounding boxes, our model can still successfully and accurately accomplish the manipulation tasks.
Figure 6: A case of detection failure. The task is to catch apples to a specified location in a normal environment.
Figure 7: A case of detection failure. The task is to catch apples to a specified location in a flittering light environment.
To test our policy’s robustness against this visual uncertainty, we introduce masks that randomly obscured bounding box input at varying ratios (0.0 to 1.0) during inference.
@@
As shown in Fig. 8, the results reveal that the policy maintains a high success rate and generalization capability when the mask ratio is below 0.4. As the mask ratio exceeds 0.4, the success rate significantly declines, leading to consistent policy failures. Coincidentally, the random mask ratio we add in the student policy training phase is 0.3. This trend reveals that our approach focusing on policy generalizability can resist visual uncertainty up to 40% of the masking ratio. That is, even if 40% of the visual information is lost, the policy still has good spatial generalizability. We fitted the data points with a Gaussian curve to represent this trend, defined by the equation:
@@
This experiment validates that our policy can withstand moderate visual uncertainties during inference, maintaining robust performance. This shows that our focus on policy generalization can easily cope with situations where the visual model has large uncertainty, leading to more robust manipulation. However, the steep decline in success rate for higher mask ratios highlights the importance of reliable visual input for effective spatial reasoning and task execution.
As to whether the random mask in the training phase gives some increase in the success rate of our policy, please refer to Sec. B.3.
B.3 Random Mask during Student Policy Training Affects Performance.
The uncertainty of visual detection in the real world is mentioned in the previous subsection, in order to demonstrate that random masking during training does favor the performance of the student policy, Specifically, we compare the success rates of policies trained with and without a random mask ratio of 0.3 across five different spatial ranges. During training, we randomly mask out 0 to 30% of the bounding box information in each trajectory. This encourages the policy to adapt to scenarios where bounding box detection is incomplete or inaccurate. Table 3 demonstrated that in all range of spaces, the policy trained with random masks shows outstanding performance compared with one without masks. In addition, the policy shows better robustness in larger spatial ranges, where the variability and uncertainty of detection are more pronounced, highlighting the effectiveness of random masking in improving spatial adaptability and generalization.
Table 3: Comparison of Success Rates With and Without Random Mask
Spatial Range Random Mask Ratio = 0.3 Without Random Mask
Fix Point 100.0% 100.0%
$5~{}\mathrm{cm}\times 5~{}\mathrm{cm}\times 5~{}\mathrm{cm}$ 83.33% 79.17%
$10~{}\mathrm{cm}\times 10~{}\mathrm{cm}\times 10~{}\mathrm{cm}$ 97.92% 64.58%
$20~{}\mathrm{cm}\times 20~{}\mathrm{cm}\times 20~{}\mathrm{cm}$ 87.5% 70.83%
Full Space ($41~{}\mathrm{cm}\times 30~{}\mathrm{cm}\times 28~{}\mathrm{cm}$) 87.5% 68.75%
B.4 Random Noise in Bounding Boxes during Inference.
To further examine the policy’s robustness against visual uncertainty, we introduce random noise to the bounding boxes during inference. We tested noise ratios at increments of 0.01, 0.02, 0.05, 0.1, 0.2, and 0.5. Fig. 9 illustrates the success rate as a function of noise ratio, where a monotonic decay fit highlights the policy’s performance trend. The results show that the policy maintains a high success rate at low noise ratios but experienced a significant decrease as noise levels increased. Since our normalized bounding box value ranges from 0 to 1, the policy still achieves a high success rate with a visual detection error of 0.05, demonstrating the robustness of the policy generalization to cope with visual uncertainty and visual sim2real gap.